# Distributed Maximum Flow in Planar Graphs

### v3

@@ 4 Weighted Girth via Minor Aggregations
We remark that Ghaffari and Parter [17] showed how to solve specific aggregate functions on $G^{*}$ using the graph $\hat{G}$ (aggregations on each face of $G$ and sub-tree sums on $G^{*}$). For our purposes, we need something more general, so we show how to perform general part-wise aggregations on $G^{*}$. In particular, we need to perform aggregations that take into consideration the outgoing edges of each part $S_{i}$, something that was not done in [17]. This specific task required our small modification to $\hat{G}$ compared to the one defined in [17].
To solve the PA problem on $G^{*}$ we exploit the structure of $\hat{G}$. We prove that a PA problem in the dual graph $G^{*}$ can be translated to a corresponding PA problem in $\hat{G}$. In the graph $\hat{G}$ there is a disjoint cycle representing each face of $G$, where two such cycles are connected if there is a dual edge between the corresponding faces. Hence, if we take a partition $\{S_{1},\ldots,S_{N}\}$ of the dual nodes, such that each $G[S_{i}]$ is connected in the dual graph, we can convert it to a partition of the vertices of $\hat{G}$, where each dual node $f\in S_{i}$ is replaced by the vertices of the cycle corresponding to $f$ in $\hat{G}$. Since cycles corresponding to different faces are vertex-disjoint in $\hat{G}$ we indeed get a partition.
In addition, by the properties of $\hat{G}$, it is a planar graph with diameter $O(D)$, and we can simulate $\mathsf{CONGEST}$ algorithms efficiently on $\hat{G}$. Based on this we can solve the PA problem in $\tilde{O}(D)$ rounds in $\hat{G}$, which leads to solving the PA problem on the dual graph $G^{*}$ in $\tilde{O}(D)$ rounds. For full details see Appendix B. We also prove that we can simulate an extended version of the minor-aggregation model defined in [18] that allows adding $\tilde{O}(1)$ virtual nodes to the network, which is useful for our applications.
Dealing with parallel edges in $G^{*}$. Recall the the dual graph $G^{*}$ can be a multi-graph, where for some of our applications it is useful to think about it as a simple graph. For example, when computing shortest paths we would like to keep only the edge of minimum weight connecting two dual nodes, and when computing a minimum cut we can replace multiple parallel edges by one edge with the sum of weights. Note that if $G^{*}$ was the network of communication it was trivial to replace multiple parallel edges by one edge locally, but in our case each edge adjacent to a face $f$ can be known by a different vertex of $f$. While we can compute an aggregate operator of all the edges adjacent to $f$ efficently, if a face $f$ has many different neighboring faces $g$, we will need to compute an aggregate operator for each pair of neighboring faces which is too expensive. To overcome it, we use the low arboricity of planar graphs to compute a low out-degree orientation of the edges of the dual graph. This orientation guarantees that from each face $f$ there are outgoing edges only to $O(1)$ other faces $g$, hence we can allow to compute an aggregate function on the edges between each such pair efficiently. The algorithm for computing the orientation is based on solving a series of $\tilde{O}(1)$ part-wise aggregation problems. For details see Appendix B.
Applications. By simulating minor-aggregation algorithms on the dual graph we can compute the weighted girth in $\tilde{O}(D)$ rounds by simulating a minimum weight cut algorithm on the dual graph and exploiting the cycle-cut duality of planar graphs (3.1). This approach allows us to find the weight of the minimum weight cycle, but we show that we can extend it to find also the edges of the cycle, for details see Appendix B. Another application is a $(1-\epsilon)$-approximation for maximum $st$-flow when the graph is undirected and $s,t$ lie on the same face with respect to the given planar embedding. Here, we simulate an approximate shortest path algorithm on a virtual graph obtained from the dual graph $G^{*}$ by slight changes, for details see Section 6.

### v4

@@ 4.2.1 PA on $G^{*}$ and Basic Model Simulation
Using aggregations on $\hat{G}$, aggregations on $G^{*}$ were previously done by Ghaffari and Parter [17] to solve the specific problem of computing a primal balanced separator, thus, their aggregation tasks were problem-specific (aggregations on faces of $G$ and sub-tree sums on $G^{*}$). For our purposes, we need something more general, so we extend their method to perform general part-wise aggregations on $G^{*}$. In particular, we need to perform aggregations that take into consideration the outgoing edges of each part $S_{i}$ (i.e., edges that have one endpoint in $S_{i}$ and the other endpoint in $S_{j}\neq S_{i}$), something that was not done in [17]. This specific task required our small modification to $\hat{G}$ compared to the one defined by Ghaffari and Parter.
Lemma 4.9 (Part-Wise Aggregations on $G^{*}$).
Let $P=\{S_{1},S_{2},\ldots,S_{N}\}$ be a partition of $V(G^{*})$, such that, (1) all $G^{*}[S_{i}]$ are connected, (2) each vertex on a face $\hat{f}$ of $\hat{G}$ that correspond to a node $f\in S_{i}$, knows the (same) ID for $S_{i}$, (3) each input of a node $f$ is known to the leader vertex of $\hat{f}$ in $\hat{G}$ and (4) each input of an edge $e^{*}$ (if any) is known to the two endpoints in $\hat{G}$ of the edge $\hat{e}\in E_{C}$ that maps to $e^{*}$. Then, any aggregate operator over $S_{i}$ nodes or edges can be computed simultaneously in $\tilde{O}(D)$ $\mathsf{CONGEST}$ rounds, such that, the output value of each $S_{i}$ is known to all vertices of $G$ that lie on the faces whose corresponding dual node is in $S_{i}$.
Proof.
The computation is done via aggregations inside $\hat{G}$, in particular, a partition of $G^{*}$ induces a partition on $\hat{G}$. Notice that $\hat{G}$ is planar, is built in $O(1)$ rounds, of diameter $O(D)$ and allows $\mathsf{CONGEST}$ simulation in $G$ with a constant blow-up factor in the round complexity (Properties 1, 2 and 3 of $\hat{G}$). Thus, $\hat{G}$ admits $\tilde{O}(D)$ quality low-congestion shortcuts in $\tilde{O}(D)$ rounds and the PA problem is solved in $\tilde{O}(D)$ rounds on it (Lemma 4.5 and Corollary 4.6).
A partition $P$ of $G^{*}$ induces a partition $\hat{P}$ on $\hat{G}$ as follows. A node $f$ of $G^{*}$ maps to a face $\hat{f}$ of $\hat{G}$, the part $\hat{S}_{i}\in\hat{P}$ is defined to contain all vertices on faces $\hat{f}$ of $\hat{G}$ that map to nodes $f$ of $G^{*}$ in $S_{i}$. All $\hat{S}_{i}$ are connected and vertex-disjoint, as each two distinct $f,g\in G^{*}$ map to distinct vertex- and edge-disjoint faces $\hat{f},\hat{g}$ of $\hat{G}$, in addition, $\hat{f},\hat{g}$ are connected by an $E_{C}$ edge for each edge in $G^{*}$ connecting $f,g$. Notice that star-center vertices ($V_{S}$) are not in any part $\hat{S_{i}}$, so we define each to be its own part. Hence, one can see that $\hat{P}=(\cup_{i\in[N]}\{\hat{S_{i}}\})\bigcup(\cup_{v\in V_{S}}\{v\})$ is a partition of $\hat{G}$ into connected vertex-disjoint subgraphs. Thus, to compute part-wise aggregation on $P$, we compute part-wise aggregations on $\hat{P}$, in which, vertices of $V_{S}$ do not participate as an input to the part-wise aggregation task and are used exclusively for communication, henceforth, we do not refer to these vertices.
Initially, all vertices in a part $\hat{S}_{i}$ know the same ID of $\hat{S}_{i}$ (equivalently, ID of $S_{i}$). Thus, $\tilde{O}(D)$ quality low-congestion shortcuts on $\hat{G}$ are constructed for the partition $\hat{P}$ in $\tilde{O}(D)$ rounds as the assumptions for constructing low-congestion shortcuts are met, and because $\hat{G}$ admits such shortcuts as mentioned above.
Recall that in the PA task, each vertex $v$ in a part of the partition has an $\tilde{O}(1)$-bit input, and the aggregate operator is computed over all inputs of vertices which are in the same part. When computing a PA task over nodes of $G^{*}$, the input of $f\in G^{*}$ is required to be known to some vertex $\hat{v}\in\hat{G}$ (w.l.o.g. the leader) on the face of $\hat{G}$ that maps to $f$ (Property 4 of $\hat{G}$). Then, aggregations on all $\hat{S}_{i}$ shall simulate aggregations on $S_{i}$, where only the input of the leader of $\hat{f}$ is considered as an input of the dual node $f$. That is, the aggregation task considers input from leaders only, and all other vertices of $\hat{G}$ participate in the algorithm just for the sake of communication (their input is an identity element).
If the PA task is over inside-part edges or over outgoing edges, then, each vertex $v$ of $\hat{G}$ that is incident to an edge $\hat{e}=\{u,v\}\in E_{C}$ (recall, edges of $E_{C}$ map 1-1 to $E(G^{*})$ by Property 5 of $\hat{G}$)), knows if $\hat{e}$’s input should or should not be considered in the aggregation - as $v$ knows the ID of the part $\hat{S}_{i}$ that contains it, and know the ID of $\hat{S}_{j}$ that contains $u$ in one communication round. Then, aggregations on all $\hat{S}_{i}$ simulate aggregations on $S_{i}$, where any $v\in\hat{G}$ that participates in the aggregation considers some of its edges as an input depending on the task and on whether they are inside-part edges or not. Vertices that have no incident input edge, participate in the algorithm just for the sake of communication (their input is an identity element). This PA task on $\hat{G}$ is computed within $\tilde{O}(D)$ rounds as mentioned at the beginning of the proof.
Now, having part-wise aggregations on $G^{*}$, we show a sketch of simulating of the minor-aggregation model in $\mathsf{CONGEST}$ via PA (as in [43]), where the input graph is the same as the network of communication. Then describe the differences with the similar proof for simulation on $G^{*}$. In the simulation, a vertex of $G$ simulates itself, and the two endpoints of an edge $e$ simulate $e$. The simulation consists of three main steps: (1) A few ($\tilde{O}(1)$) part-wise aggregation tasks are executed to elect a leader for each super node, defining the nodes of the obtained minor, (2) A part-wise aggregation task is executed to simulate consensus, where a part is $G[s]$ ($s$ is some super-node), and finally (3) The endpoints of each edge exchange their parts’ consensus values and one more part-wise aggregation task is performed to compute the aggregation values. Lemmas 4.8 and 4.9 pave the way for the following theorem.
Theorem 4.10 (Basic minor-aggregation model for $G^{*}$).
Proof.
Let $s=\{f_{1},f_{2},\ldots\}$ be a super-node of $G^{*}$ ($f_{i}\in V(G^{*})$), the vertices of $\hat{G}$ that map to $s$ are the vertices that lie on the faces $\hat{f}_{i}$ of $\hat{G}$, such that, $\hat{f}_{i}$ maps to the node $f_{i}$. Then, the computational power of a super-node of $G^{*}$ is simulated by all vertices of $\hat{G}$ that map to it. The computational power of an edge in (a minor of) $G^{*}$ shall be simulated by the endpoints of its corresponding $E_{C}$ edge in $\hat{G}$ (Property 5). Often, we will use Lemma 4.9 as a black-box for PA on $G^{*}$ in order to implement the simulation.
  1. 1.
    Electing a leader (and contraction). Recall, in the contraction step, each edge has a value of zero or one, and considered to be contracted if and only if its value is one. To implement that, we have mainly two computational tasks. First, merge clusters of nodes of $G^{*}$ over their outgoing edges of value one (clusters eventually define super-nodes of $G^{*}$). Second, assign each cluster an ID that is known to all vertices of $\hat{G}$ that simulate it. Initially, a cluster is a single node of $G^{*}$, then, clusters are grown to match the set of super-nodes of (the minor of) $G^{*}$.
    Implementing that, (1) connected components of $\hat{G}[E_{R}]$ are detected (those are faces of $\hat{G}$ that map to nodes $v\in G^{*}$), where the ID of a node of $G^{*}$ is the minimal ID of a vertex in $\hat{G}$ that maps to it, this is done in $\tilde{O}(D)$ rounds by property 4 of $\hat{G}$. Now each vertex of $\hat{G}$ knows in which of these connected components it is (defines a partition on $G^{*}$ where each part is a single node) defining each node as a cluster. Notice that now Lemma 4.9 (PA in $G^{*}$) is applicable, due to the fact that all vertices of $\hat{G}$ that correspond to a node $f\in G^{*}$ in a cluster, know the same ID for that cluster. Then, (2) we grow clusters. The merging process follows Boruvka’s classic scheme, where the process breaks down into $O(\log n)$ star-shaped merges, the classic implementation is randomized and similar to that of [14, 17], however, one can implement this merging process deterministically with the derandomization of [18]. Namely, the PA task on $G^{*}$ here is when each cluster defines a part in a partition of $V(G^{*})$, and the aggregate operator is computed over the outgoing edges of each cluster that chose a value of 1 in the contraction step, deciding the merges; the operator computed is over labels of neighboring clusters incident to those edges, where each cluster is labeled as a joiner or a receiver, such that, joiners suggest merges and receivers accept them (i.e., receivers are the star centers in the star-shaped merges). Randomly, the labels are decided by a fair coin flip that the leader of the cluster tosses and broadcasts to the whole cluster. Deterministically, the classic algorithm for 3-coloring by Cole and Vishkin [3] is used (E.g. see [18] for more details). At most $\tilde{O}(1)$ merges are repeatedly done until clusters match super-nodes of $G^{*}$. In each phase of merges, all nodes maintain the minimum ID of a node $v$ (of $G^{*}$) in the cluster. The vertex of $\hat{G}$ that has that exact ID is considered to be the leader of vertices in $\hat{G}$ that map to the super-nodes represented by the cluster. Notice, there is such a vertex in $\hat{G}$ that has that exact ID since IDs of nodes of $G^{*}$ are IDs of vertices in $\hat{G}$ that map to them. In addition, at the end of each merging phase Lemma 4.9 (PA on $G^{*}$) is applicable, thus, it can be applied repeatedly $\tilde{O}(1)$ times, once for each merging phase. By Lemma 4.9, computing $\tilde{O}(1)$ PA tasks on $G^{*}$ is done in $\tilde{O}(D)$ rounds on $G$.
    The above implements the contraction step, after which, all vertices of $\hat{G}$ that simulate the same super node of $G^{*}$ have a cluster-unique shared leader and ID.
  2. 2.
    Consensus. After assigning IDs to clusters, as said, each vertex of $\hat{G}$ knows the ID of the cluster that it maps to, which is used as the part ID in the part-wise aggregation tasks on $G^{*}$. Regarding input, the input is (perhaps unique) per node of $G^{*}$, which elected leaders in $\hat{G}$ know (those leaders were elected in the first phase of cluster merging in the previous step). Hence, Lemma 4.9 is applicable and is used to implement the consensus step in $\tilde{O}(D)$ rounds on $G$.
  3. 3.
    Aggregation. After performing the consensus step, each endpoint $u$ in $\hat{G}$ of an edge in $E_{C}$ informs the other endpoint $v$ with the consensus value of the cluster that $u$ maps to. Then, an endpoint $v\in\hat{G}$ of an edge $e\in E_{C}$ that maps to the edge $e^{*}=(a_{1},a_{2})$ (where $a_{1},a_{2}$ are distinct super nodes of $G^{*}$), does the following. If $v$ is in the cluster (that maps to) $a_{i}$, $v$ thinks of the edge $e$ as if it was assigned the value/weight of $z_{e,a_{i}}$, and an identity element otherwise. Then, the aggregate operator is computed over the outgoing edges of super-nodes. Again, the aggregation is done by Lemma 4.9 in $\tilde{O}(D)$ rounds on $G$. Finally, the two endpoints of each edge know the output of the aggregate operator, thus, they can choose zero or one for that edge, and we are ready for the next round of minor-aggregation.
Notice that after the simulation finishes, each vertex in a cluster knows the output value of its cluster’s corresponding super-node. Hence, by definition of $\hat{G}$ (see Property 1 of $\hat{G}$), for each $f^{*}\in G^{*}$ its output is known to all vertices on the corresponding face $f$ of $G$, and the value of an edge $e^{*}$ that maps to $e\in\hat{G}$ is known to $e$’s endpoints in $\hat{G}$ as specified in the statement of Lemma 4.9. ∎
4.2.2 Extended Model Simulation
We define now the extended minor-aggregation model as presented in [18] and show that it can be simulated in $G^{*}$.
Definition 4.11 (Extended minor-aggregation model [18]).
This model extends the minor-aggregation model (Definition 4.7) to work on virtual graphs that have a small number of virtual nodes, without strongly affecting the round complexity. In particular, (1) arbitrarily connected $\tilde{O}(1)$ virtual nodes can be added to the network, and (2) $\tilde{O}(1)$ nodes of the network, each can be replaced with an arbitrarily connected virtual node.
Lemma 4.12 (Lemma 15 and Theorem 14 of [18]).
Lemma 4.13 (Lemma 15 and Theorem 14 of [18]).
Our proof would make a black-box use of Lemma 4.13, however, we provide a sketch of the proof of [18] just to make things a bit clearer (for the full proof see Theorem 14 of [18]). (1 - Contraction) Contract non-virtual edges that shall be contracted, then, in $O(\beta)$ rounds, each super node shall learn the IDs of virtual-nodes it is connected to, via computing an aggregate over its incident virtual edges. Then, the super-nodes’ IDs get computed, (2 - Consensus) At first, all super-nodes that do not contain any virtual node perform their consensus step, then, we iterate over all virtual nodes and compute the consensus of the super-node they are contained in, such that, all nodes of the graph know the output value of each super node. Finally (3 - Aggregation) A real edge simulates itself, a virtual edge that has one non-virtual endpoint node is simulated by that endpoint and a virtual edge that the two of its endpoints are virtual is simulated by all nodes. The $z$ value of each edge is computed and an aggregation is performed as in the consensus step.
Simulating the extended minor-aggregation model to run on $G^{*}$ is a direct corollary of Lemma 4.12, Lemma 4.13, and Theorem 4.10. Let $G_{virt}^{*}$ be a virtual graph that is obtained from $G^{*}$ by adding and replacing nodes with up to $\beta$ arbitrarily connected virtual nodes, then,
Theorem 4.14.
Notice that adding virtual nodes and edges to $G^{*}$ might violate planarity, but that is not a problem, as these virtual nodes and edges are simulated by (on top of) the same underlying network $\hat{G}$ that is used to simulate the basic model, without making any changes on it. In particular, this extended model, is simulated by the basic model unconditionally, i.e., Lemma 4.13 and 4.12 constitute a universal algorithm that runs in the basic model for simulating the extended model, so we simply apply that algorithm on $G^{*}$ using Theorem 4.10.
4.2.3 Dealing with Parallel Edges in $G^{*}$
We now address the problem of $G^{*}$ being a (multi) non-simple graph, i.e. has self-loops and parallel edges. This might be a problem due to the fact that many algorithms assume their input graph is simple. We provide a procedure for deactivating parallel edges of $G^{*}$: Consider two nodes $u,v$ of $G^{*}$ that are connected by multiple edges $e_{1},e_{2},\ldots,e_{m}$. We deactivate all these edges except for one (called the active edge). If edges are assigned a weight $w(\cdot)$, the active edge gets assigned a new weight $\otimes_{i}w(e_{i})$, where $\otimes$ is a predefined aggregation operator. E.g., $\otimes$ is the summation operator if one is interested in computing a minimum cut, and the minimum operator if one is interested in computing shortest paths.
One cannot straightforwardly use part-wise aggregations on $G^{*}$ in order to do the above, since a node of $G^{*}$ might be connected to many nodes, and the required task is an aggregation for each neighbor rather than one aggregation over all incident edges, this may lead to high congestion if all nodes do those aggregations naively and simultaneously. We use the the fact that the underlying simple graph of $G^{*}$  [^7] I.e. $G^{*}$ after the removal of self-loops and ”collapsing” parallel edges to a single edge. has a small arboricity of $\alpha=3$ (the arboricity of a graph is the minimal number of forests into which its edges can be partitioned) to obtain a minor-aggregation procedure that requires $\tilde{O}(\alpha)$ minor-aggregate rounds for orienting $G^{*}$ edges, so that in the underlying simple graph of $G^{*}$ each node has at most a constant out-degree, which we utilize for load balancing the aggregation tasks. In the original $G^{*}$ (the multi graph) this means that each node would have outgoing edges to a constant number of neighbors, rather than a constant number of outgoing edges. I.e., we allow many edges to be oriented outwards a node as long as they connect it to a constant number of neighbors. We then apply this procedure on $G^{*}$ in a black-box manner using Theorem 4.14, resulting with an $\tilde{O}(D)$ $\mathsf{CONGEST}$ algorithm on $G$ simulating the procedure on $G^{*}$. We henceforth assume that $G^{*}$ is a simple graph. Formally, we obtain the following lemma.
Lemma 4.15 (Deactivating parallel edges and self-loops).
Let $\alpha$ be the arboricity of the input graph, we implement an algorithm of [1] in the minor-aggregation model that runs in $\tilde{O}(\alpha)$ minor-aggregation rounds and produces such an orientation. Since the arboricity of the underlying simple graph of $G^{*}$ is three, we get our constant out-degree orientation.
Proof.
First, we omit self loops, each edge whose endpoints have the same ID, deactivates itself, this is done in only one minor-aggregate round. Next, we describe [1]’s algorithm and then describe how to implement each step in the minor-aggregation model. At the start of the algorithm, all nodes are white. For each node, there is a time where it becomes black, the algorithm terminates when all nodes turn black. The algorithm produces a partition of the node set into $\ell=\lceil 2\log n\rceil$ parts, $H_{1},\ldots,H_{\ell}$. Upon termination, each node knows the part $H_{i}$ that contains it. Moreover, the following orientation of edges, results with an orientation with at most $O(\alpha)$ outgoing edges for each nodes (counting parallel edges between two nodes as one, that is, in the underlying simple graph). The orientation is as follows. An edge $e=(u,v)$ where $u\in H_{i}$ and $v\in H_{j}$ s.t. $i<j$, gets oriented towards $v$. Otherwise, if both $u,v$ belong to the same part, $e$ is oriented towards the endpoint with a greater ID.
  1. 1.
  2. 2.
    If the super-node $v$ has counted $3\alpha$ neighbors, via a consensus step, this information ($v$ shall turn black) is broadcasted to all nodes inside $v$. I.e., edges incident to $v$ learn this information, and $v$’s neighbors learn that $v$ is black and belongs to $H_{i}$.
After the algorithm terminates, each edge knows to orient itself as was described earlier. Each node now has at most $\alpha$ outgoing edges (counting parallel edges as one). This procedure terminates in $\tilde{O}(\alpha)$ rounds. Note, the orientation of an edge can be decided locally by the two vertices in the communication graph which simulate that edge (in our case, these are vertices of $\hat{G}$).
Deactivate parallel edges. Assume edges are assigned unique $\tilde{O}(1)$-bit IDs in $\tilde{O}(1)$ rounds. Having self-loops deactivated, we now use the orientation to deactivate parallel edges. Each node deactivates its outgoing parallel edges in $O(\alpha)$ rounds. First, enumerate outgoing edges from $1$ to $3\alpha$, where all edges that connect $v$ to the same neighbor have the same number, this is done in $O(\alpha)$ rounds, in an iterative way similar to that of counting neighbors. Then, for all $i\in[3\alpha]$, for edges with number $i$: (1) An aggregation step over their weights is computed, (2) An edge is chosen to be the active edge by another aggregation step (say the minimum ID one) and other edges deactivate themselves. This process terminates in $O(\alpha)$ rounds. ∎
4.3 Weighted Girth
In this subsection we show how to compute the weighted girth (i.e. the minimum-weight cycle) of an undirected weighted planar graph $G$. We do this by exploiting the result of Section 4.2 in order to simulate a minor-aggregate weighted minimum-cut algorithm on $G^{*}$, which (by 3.1) gives the weighted girth of $G$.
[…]
Theorem 4.16 (Minor-aggregate exact min-cut, Theorems 12, 18 and 40 of [18]).
Let $G$ be an $\exp(\tilde{O}(1))$ edge-weighted input graph. There is a randomized $\tilde{O}(1)$-round minor aggregation algorithm that computes the weight of the minimum weight cut $C$ of $G$ w.h.p.. Upon termination, the value of $C$ is known to all vertices.
Notice, applying LABEL:{th:_GZ_min_cut} above on $G^{*}$ already allows us to find the weighted girth of our planar network in $\tilde{O}(D)$ rounds. In [18]’s work, the focus was on finding the weight of the minimum cut, but it can be easily extended to find also the edges of the minimum cut. Let $T$ be a spanning tree of $G$ and $C=(S,V(G)\setminus S)$ be a cut of $G$, then, $C$ is, respectively, (one-) two-respecting of $T$ if and only if there are exactly (one) two edges of $T$ that cross the cut. Given such a tree, where its edges that cross the cut are known, in $O(1)$ minor aggregation rounds, one can mark all edges of the cut.
Lemma 4.17 (Mark cut edges).
Let $G$ be any network, $T$ a spanning tree of $G$ and $C$ any cut that one- (two-) respects $T$, defined by edges $e_{1},e_{2}\in T$ (possibly $e_{1}=e_{2}$). Given $G$, assuming that edges of $T$ and the edges $e_{1},e_{2}$ know themselves, there is an $O(1)$ deterministic minor-aggregation algorithm that marks all $C$ edges in $G$ (i.e., upon termination, each edge $e\in C$ knows itself as such).
Proof.
Consider the tree $T$, we start by (a contraction step) contracting all $T$ edges except for $e_{1},e_{2}$, which results in a minor $G^{\prime}$ of $G$ with at most three super nodes that are connected by $e_{1},e_{2}$ and possibly other (parallel) edges. Next, each super-node shall compute its cost (defined as the number of edges from $e_{1},e_{2}$ that are incident to it). This is done by an aggregation step, each edge in $G^{\prime}$ chooses a weight of zero, except for $e_{1},e_{2}$, each chooses a weight of one, and an aggregation step is performed with the sum operator, resulting with the sum of outgoing edges for each super-node, which is its cost. Now, the ID of the maximum-cost node is computed and broadcast to all the graph - by contracting all the graph into a single super-node $s$, initiating the input of each node $v$ to a tuple that contains the ID and the cost of the super-node $s^{\prime}\in G^{\prime}$ s.t. $v\in s^{\prime}$, then, we compute the maximum cost by a consensus step (breaking ties arbitrarily). Finally, we return to consider $G^{\prime}$ as our graph (we undo the last contraction). Notice, edges of $G^{\prime}$ know the ID of the maximum-cost node, so, they mark themselves if only if they are incident to it.
The round complexity is clear. The correctness is as follows. $G^{\prime}$ is a graph that is composed of two or three super-nodes, each super-node corresponds to a connected component of $T\setminus C$, each induces a connected subgraph of $G$, such that, one of which is incident to all edges $e_{1},e_{2}$ in $G$, that subgraph defines one side of the cut (denoted $S$) and the other (at most two) subgraphs define the other side of the cut. Since $S$ is incident to $e_{1},e_{2}$ it has the highest cost, thus, all edges exiting $S$ in $G$ get marked, these are exactly the edges of $G$ that cross the cut, hence, all edges of $C$. ∎
The algorithm of [18] (Theorem 4.16), like many min-cut algorithms (both centralized and distributed, see [24, 25, 44, 14, 12, 11, 33, 5]) consists of three main steps: (1) pack spanning trees of $G$, (2) compute a minimum 1-respecting cut of each packed tree and (3) compute a minimum 2-respecting cut of each packed tree, the final output is the overall minimum. As said, upon termination the value of $C$ is known to all nodes, in addition, it is implied in [18] that the edges of a spanning tree $T$ of $G$ that $C$ (1-) 2-respects know themselves, in particular, edges in $T\cap C$ know themselves as such. Using Theorem 4.14 on $G^{*}$ in order to apply the exact min-cut algorithm (Theorem 4.16) and to apply the procedure for marking cut edges (Lemma 4.17), we get an undirected planar weighted girth algorithm:
See 1.7
Proof.
First we make $G^{*}$ simple by applying Lemma 4.15 for deactivating parallel edges, the (only) active edge left between any two nodes $u,v$, gets assigned a weight equal to the sum of weights of edges with the same endpoints in $G^{*}$. Then, we compute the minimum cut $C$ of the dual network $G^{*}$ by applying Theorem 4.14 to simulate [18]’s minor-aggregate exact min-cut algorithm (Theorem 4.16) on $G^{*}$. Afterwhich, we can assume that we have the min-cut $C$ value, that $C$ is (1-) 2-respecting to some tree $T$, and that each edge in $T$ along with $T$’s edges $e_{1},e_{2}$ that cross the cut know themselves ($e_{1}=e_{2}$ in case $C$ is 1-respecting). Now we use Theorem 4.14 for applying Lemma 4.17 in order to mark the edges of $C$ in $G^{*}$. By the cycle-cut duality (3.1), the weight of $C$ is the weight of $G$’s minimum cycle and the edges of $C$ constitute a minimum weight cycle in $G$. Since all procedures we use on $G^{*}$ are of $\tilde{O}(1)$ minor-aggregate round complexity, then, the $\mathsf{CONGEST}$ round complexity for simulating them over $G$ is $\tilde{O}(D)$ by Theorem 4.14. ∎
We clarify a low-level detail regarding marking the edges of $C$. Recall, the minor-aggregation model simulation on $G^{*}$ is done via the face-disjoint graph $\hat{G}$. Since the edges of $C$ are edges of $G^{*}$, each edge $e^{*}$ of them maps to exactly one edge $\hat{e}$ of $\hat{G}$ (Property 5), which itself maps to the edge $e\in G$ (the dual of $e^{*}$), that is, indeed, marking edges of $G^{*}$ marks the corresponding edges of $G$. Note however, edges $\hat{e}\in\hat{G}$ that map to marked edges might not form a cut nor a cycle in $\hat{G}$ but that is fine as it is only a communication tool and not our input graph for the problem.